### v2

# On Variational Perspectives To The Graph Isomorphism Problem

Abstract
We consider a quadratic unconstrained binary optimization (QUBO) formulation of the graph isomorphism problem from a variational quantum algorithmic perspective. By treating it using the Quantum Approximate Optimization Algorithm (QAOA) and the Variational Quantum Eigensolver (VQE), this study presents results for each and highlights variations therein for graphs of four and five nodes. For isomorphic graphs having an equal number of nodes and edges, we demonstrate clustering in the energy landscape for the QAOA. This trend found in the QAOA study was further reinforced by studying the ground state energy reduction using VQEs. Furthermore, we examine the trend under which isomorphic pairs of graphs vary in the ground state energies, with varying edges and nodes.
preprint: AIP/123-QED
I Introduction
The graph isomorphism problem holds a special place in theoretical computer science because it sits in a gray area of computational complexity. While not known to be NP-complete, it is also not proven to be solvable in polynomial time, and Babai’s breakthrough result Babai (2016) established a quasi-polynomial time classical algorithm — a landmark in complexity theory. Beyond its foundational role, the problem appears in diverse areas such as molecular similarity in chemistry, protein interaction networks in biology, and structural equivalence in social networks Doga et al. (2024); Mohtashim (2025). These applications underscore both the theoretical depth and the practical importance of efficient quantum approaches.
Hybrid quantum-classical algorithms have emerged as a leading strategy to obtain quantum advantage in Noisy Intermediate Scale Quantum (NISQ) devices Kyriienko, Umeano, and Holmes (2025); Cerezo et al. (2025); Mohtashim et al. (2025); Mohtashim, Mahatabuddin, and Jabbar (2023); Brokowski et al. (2022); Cerezo et al. (2021); Preskill (2018). In particular, variational quantum algorithms (VQAs), which provide a general framework for tackling important classes of problems, involve quantum circuits whose parameters are are optimized by classical optimisation techniques Magann et al. (2021). Typically in VQAs, like the Variational Quantum Eigensolver (VQE) Peruzzo et al. (2014) and Quantum Approximate Optimization Algorithm (QAOA) Farhi, Goldstone, and Gutmann (2014), the goal is to find a quantum state which minimizes the energy of a problem Hamiltonian. Earlier, quantum annealing (QA) had enabled quantum-enhanced information processing for roughly the same areas of application as VQAs Lucas (2014). The key difference between VQAs and QA is that the former involves a gate-based computation, and the latter is quantum-fluctuation-based computation exploiting quantum mechanical effects such as tunneling and entanglement. Quantum Annealers are relatively easier to build and have significantly more qubits than gate-based quantum devices Venegas-Andraca et al. (2018).
What underpins quantum annealing is the adiabatic theorem in physics, which involves slowly evolving the ground state of a physical system from the problem Hamiltonian of the system, and is what enables finding the ground state energies relevant to QUBO problem formulations Santoro and Tosatti (2006); Lewis and Glover (2017); Domino et al. (2021). Most prior quantum information efforts to tackle the graph isomorphism problem have focused on quantum annealing approaches. While annealers are effective for specific instances, they face limitations in scalability, connectivity, and flexibility. In contrast, VQAs provide greater adaptability to different problem formulations on account of relying on a gate-based circuit model. They are furthermore empowered by the hybrid quantum-classical optimization loop, making them more robust for implementations on noisy intermediate-scale quantum (NISQ) devices. By considering VQAs for tackling graph isomorphism, this work explores whether the variational paradigm offers not only practical performance on small graph instances, but also a potential pathway to scalable quantum heuristics for one of the most enigmatic problems in complexity theory.
Prior QUBO formulations for the Graph Isomorphism Problem have been provided by LucasLucas (2014), Kenneth Zick, Shehab, and French (2015), Calude Calude, Dinneen, and Hua (2017) etc. Kenneth and Calude performed their QUBO formulation experiments on D-Wave’s Quantum Annealer machines. It has been conjectured that the lowest number of logical qubits required for the general Graph Isomorphism Problem for two graphs of order n is $n^{2}$. Apart from this, leveraging QUBO for a problems such as Boolean satisfiability Farhi et al. (2000), binary clustering Bauckhage et al. (2017), classifier training Pudenz and Lidar (2013), Travelling Salesman Problem Salehi, Glos, and Miszczak (2021) vehicle routing Borowski et al. (2020), railway dispatching problem Domino et al. (2021) has been explored. This work presents a first proof-of-principle approach to solving the Graph Isomorphism Problem using the circuit model of quantum computation through variational means in a NISQ paradigm.
II Background
The Graph Isomorphism Problem is the computational problem of determining whether two finite graphs are structurally identical or isomorphic. An isomorphism of a graph $G=(V,E)$ to a graph $H=(W,F)$ is a one-to-one, bijective mapping from the vertex set of the first graph $V$ to the vertex set of the second graph $W$ that preserves adjacency and non-adjacency. The exact complexity of the problem has eluded researchers ever since its conception for more than five decades. The problem is not known to be solvable in polynomial time generally, but interestingly, for many special classes of graphs, the problem becomes amenable to polynomial time solutions Ladner (1975). If the graphs have different sizes or orders, then they cannot be isomorphic, and these cases can be decided quickly. The current best known classical algorithm for the problem has a time complexity of $\exp\left((\log n)^{O(1)}\right)$ which is quasi-polynomial. Babai (2016) The quantum computational complexity of the problem is unknown Shehab and Jr. (2017). Previous attempts using quantum algorithms in the light of Hidden Subgroup Problem, using Quantum Fourier Sampling has been shown to fail Hallgren et al. (2010); Grigni et al. (2001); Jozsa (1998).
II.1 The Quantum Approximate Optimization Algorithm
QAOA, introduced by Farhi et.al Farhi, Goldstone, and Gutmann (2014), finds an approximation to the ground state of a problem Hamiltonian by constructing a specific variational ansatz through first order Suzuki-Trotter decomposition Lancaster and Allen (2025); Alvarez et al. (2024); Lee (2023) approximating adiabatic evolution. The operators $\textrm{exp}(-i\beta H_{\text{m}})$ and $\textrm{exp}(-i\alpha H_{\text{o}})$ are applied in alternation resulting in the state
$\lvert\mathbf{\alpha},\mathbf{\beta}\rangle=\prod_{k=1}^{L}\textrm{exp}(-i\beta_{k}H_{\text{m}})\textrm{exp}(-i\alpha_{k}H_{\text{o}})\lvert+\rangle^{\otimes n},$ (1)
where $H_{\text{o}}$ is the objective and $H_{\text{m}}=-\sum_{j}X_{j}$ is the mixer Hamiltonian. Under fixed number of layers (i.e. $L$), the algorithm requires $2L$ parameters. The expectation value $\langle\alpha,\beta|H|\alpha,\beta\rangle$, of state $\lvert\alpha,\beta\rangle$ is approximated through measuring the state in the computational basis. The parameters $\alpha$, $\beta$ are updated using classical optimization procedures so that the energy returns an optimal value. QAOA can be harnessed to solve problems such as Max-CutFarhi et al. (2001); Majumdar et al. (2021), Max k-Vertex CoverCook, Eidenbenz, and Bartschi (2020) and even integer factorization Anschuetz et al. (2019).
II.2 The Variational Quantum Eigensolver
The VQE uses Ritz’s variational principle to prepare an approximation to the ground state of a Hamiltonian and to estimate its corresponding energy. In this algorithm, the quantum device is tasked with preparing parameterized trial states
$\lvert\psi(\mathbf{\theta})\rangle=U(\mathbf{\theta})\lvert 0\rangle,$
where $U(\mathbf{\theta})$ is a parameterized unitary depending on a set of classical parameters $\mathbf{\theta}$. The expectation value of the Hamiltonian,
$E(\mathbf{\theta})=\langle\psi(\mathbf{\theta})|H|\psi(\mathbf{\theta})\rangle,$
is measured on the quantum hardware, and a classical optimizer is employed to iteratively update $\mathbf{\theta}$ until the minimum is reached. The variational principle ensures that every evaluated energy provides an upper bound on the true ground state energy, making the method both flexible and robust.
A key advantage of VQE over purely classical methods is its ability to efficiently explore quantum states that are exponentially hard to represent on a classical computer. This is particularly significant for problems where correlations or entanglement play a central role. Consequently, VQE has become one of the most widely studied near-term quantum algorithms, with applications spanning quantum chemistry, condensed matter physics, materials science, and combinatorial optimization Mohammad, Pivoluska, and Plesch (2024). Despite its usefulness, VQE is computationally costly as each optimization step requires repeated circuit preparation and measurement, which quickly becomes expensive as system size grows. Moreover, the choice of ansatz and optimizer strongly affects performance, and noise in near-term devices adds additional challenges Mohtashim et al. (2025); Brokowski et al. (2022). However, these costs are balanced by its unique strengths: (i) it provides a variational certificate of accuracy through Ritz’s principle, (ii) it allows flexible tailoring of ansatz circuits to exploit problem structure, and (iii) it can act as a high-precision validator of phenomena first observed in lighter, more approximate algorithms such as QAOA. In the context of the Graph Isomorphism problem, the use of VQE is particularly valuable as a complementary check. Whereas QAOA reveals clustering in energy landscapes across isomorphic graph families at relatively low circuit depth, VQE confirms the same phenomenon with greater fidelity, albeit at a higher computational cost. This duality—fast approximate detection via QAOA and rigorous confirmation via VQE—makes VQE an essential component in our variational study, demonstrating that the observed clustering effect is not an artifact of shallow circuits but rather an intrinsic property of the underlying Hamiltonian encoding of isomorphism.
III Formulation of quadratic program
The first described baseline penalty Hamiltonian for the graph isomorphism was introduced in Lucas Lucas (2014). Let us consider $G_{1}=(V_{1},E_{1})$, and $G_{2}=(V_{2},E_{2})$ graphs, where $V$ are the vertices and $E$ edges. The Hamiltonian is formulated by proposing binary variables $x_{v,i}$, which is $1$ if the vertex $v$ in $G_{2}$ gets mapped to vertex $i$ in $G_{1}$ as follows:
$H=A\sum_{v}\left(1-\sum_{i}x_{v,i}\right)^{2}+A\sum_{i}\left(1-\sum_{v}x_{v,i}\right)^{2}+B\sum_{ij\notin E_{1}}\sum_{uv\in E_{2}}x_{u,i}x_{v,j}+B\sum_{ij\in E_{1}}\sum_{uv\notin E_{2}}x_{u,i}x_{v,j}.$ (2)
In Zick et.al Shehab and Jr. (2017) a modified formulation of Eq.2 is proposed in order to achieve a simple set of coupler values (e.g. 0, 1 instead of 0, 1, 2 in Lucas) amenable to quantum annealing. An edge related penalty is applied to a coupling iff there is not a vertex mapping penalty. The modified Hamiltonian is given as follows:
$\begin{split}H=A\sum_{v}\left(1-\sum_{i}x_{v,i}\right)^{2}+A\sum_{i}\left(1-\sum_{v}x_{v,i}\right)^{2}+B\sum_{\begin{subarray}{c}ij\notin E_{1}\\ i\neq j\end{subarray}}\sum_{uv\in E_{2}}x_{u,i}x_{v,j}+B\sum_{ij\in E_{1}}\sum_{\begin{subarray}{c}uv\notin E_{2}\\ u\neq v\end{subarray}}x_{u,i}x_{v,j}.\end{split}$ (3)
In Calude Calude, Dinneen, and Hua (2017) a different methods for constructing efficient QUBO formulations for the Graph Isomorphism Problem. The objective for the formulation is given as follows:
$F(x)=H(x)+\sum_{ij\in E_{1}}P_{i,j}(x),$ (4)
where
$H(x)=\sum_{0\leq i<n}\left(1-\sum_{0\leq i^{\prime}<n}x_{i,i^{\prime}}\right)^{2}+\sum_{0\leq i^{\prime}<n}\left(1-\sum_{0\leq i<n}x_{i,i^{\prime}}\right)^{2},$ (5)
and
$\sum_{ij\in E_{1}}P_{i,j}(x)=\sum_{0\leq i^{\prime}<n}\left(x_{i,i^{\prime}}\sum_{0\leq j^{\prime}<n}x_{j,j^{\prime}}\left(1-e_{i^{\prime},j^{\prime}}\right)\right),$ (6)
$e_{i,j}=1$ if $ij\in E_{2}$ and $e_{i,j}=0$ if $ij\notin E_{2}$. This direct formulation of the quadratic function is done such that a mapping $f$ can be decoded from that values of the variables in the quadratic program using a partial function $D$, such that $D:\mathbb{Z}_{2}^{n^{2}}\mapsto\mathcal{F}$, where $\mathcal{F}$ is the set of all bijections between $V_{1}$ and $V_{2}$. The domain of $D$ consists of all vectors $\textbf{x}\in\mathbb{Z}_{2}^{n^{2}}$. Calude et. al. ensures that the function $H(x)=0$ if and only if $D$ is defined (and bijective). The component $\sum_{ij\in E_{1}}P_{i,j}(x)=0$ if and only if the mapping $f=D(\textbf{x})$ is edge invariant. Therefore the function $F(\textbf{x})=0$, for all $\textbf{x}\in\mathbb{Z}_{2}^{n^{2}}$ if and only if $f=D(\textbf{x})$ is an isomorphism.
IV Methods
The methodology for each experiment is different but was done on a common dataset unless stated otherwise. The data generator uses the NetworkX library and generates pairs of isomorphic and non-isomorphic connected graphs. The generated graphs can be found in the GitHub repository mentioned in the Data Availability section. For graphs having four nodes, the edge count remains in between four and five. On the other hand, for graphs having five nodes, all the graphs have five edges. Since the size of the QUBO matrix for a connected pair of graphs having $N$ nodes is $N\times N$, thereby utilizing $N^{2}$ qubits, the problem sizes considered for this work were kept small for NISQ era hardware simulations and computational constraints. The QUBO matrix was generated using GSGMorph If and only If (Iff) from which the quadratic programs were obtained. For each set of problem instances, the offset value for all experiments is determined empirically such that the problem yields an optimal function value of $\theta$ when the graphs are isomorphic, and the optimal function would raise a positive value when the graphs are non-isomorphic. This offset tuning is necessary because QUBO encodings can introduce constant energy shifts; setting the baseline ensures that isomorphic cases yield exactly zero cost.
IV.1 Numerical Setup for QAOA
The Quantum Approximate Optimization Algorithm (QAOA) was implemented in Qiskit using the QAOA class. The quadratic program, formulated following the approach of Calude et al. Calude, Dinneen, and Hua (2017), was first expressed in IBM’s Docplex framework, which enabled its direct translation into the corresponding problem Hamiltonian. This Hamiltonian was then supplied to the QAOA class through the minimum eigensolver interface, which evaluates the minimum eigenvalue associated with the encoded cost function. We restrict to p = 1 to remain in the shallow-circuit regime typical of NISQ demonstrations, balancing fidelity with feasibility.
IV.2 Numerical Setup for VQE
The Variational Quantum Eigensolver was implemented using the VQE class in Qiskit. The quadratic program was first expressed in IBM’s Docplex optimization framework, which served as the interface for constructing the corresponding problem Hamiltonian. For the variational ansatz, we employed the TwoLocal circuit consisting of alternating layers of single-qubit $R_{Y}$ rotations and controlled-$Z$ entangling gates, with five repetitions and a linear entanglement topology. The optimization of circuit parameters was performed using the Simultaneous Perturbation Stochastic Approximation (SPSA) optimizer, with a maximum of 600 iterations, as implemented within Qiskit’s MinimumEigenOptimizer class.
((a))
((b))
Figure 1: The energy values of the QAOA for which the resulting values of their corresponding quadratic values yield an optimal function value of zero. (a) Shows the plot of the energy values for which the values of the quadratic program for the Graph Isomorphism Problem for a four node graph yields a function value of zero. (b) Shows the plot of the energy values for which the values of the quadratic program for the Graph Isomorphism Problem for a five node graph yields a function value of zero. The boxplot shows clearly the distribution of values, the medians and standard deviations for the absolute energy values.
IV.3 Limitations
Our experiments were restricted to graphs of order four and five due to both hardware and algorithmic constraints. In the QUBO formulation, the number of logical qubits scales as $\mathcal{O}(n^{2})$ with the number of vertices $n$, making even moderately larger instances infeasible on current NISQ devices with limited qubit counts, connectivity, and coherence times. Classical simulation is likewise hindered by the exponential growth of the Hilbert space. Further limitations arose from algorithmic overhead. The VQE, even with efficient optimizers such as SPSA, is challenging to scale to larger graphs, and remains resource-intensive. For these reasons, we confined our study to small graphs in order to establish proof-of-principle results. This choice enabled us to demonstrate energy clustering for isomorphic graph families under controlled conditions, while leaving the extension to larger and more complex graphs to future work. Progress in compact QUBO encodings, problem-tailored ansätze, and error mitigation techniques will be necessary to overcome these current limitations.
V Results and Discussion
V.1 QAOA Results
The energies calculated are of the minimum eigenvalues of the converted problem Hamiltonian from the initial quadratic Docplex model using the MinimumEigenOptimizer class that takes the quadratic model and converts them into a tensor product of Pauli matrices. The corresponding minimum energy eigenvalues of the Hamiltonian for which the quadratic program yields a minimum eigenvalue of zero are given in Fig.˜1(a). Similarly, for graphs with five nodes, the energies are presented in Fig.˜1(b). Extending on the results by Szegedy Szegedy (2019), in the light of the Ising formulation Bonati (2014); Istrail (2000); Barahona (1982) of the Graph Isomorphism Problem, solved using the QAOA, we observe that the energies for graphs having the same node count have roughly the same energies, as is evident from the graphs. Moreover, the number of edges have roughly the same energies, with the energies values being distributed in a space having a relatively small value of standard deviations.
Szegedy’s result Szegedy (2019) states that for a MAX-CUT formulation of QAOA, two isomorphic graphs will yield the same energy values given the same parameters for the QAOA quantum circuit. Our results support this by showing that families of isomorphic graphs have energies within a very short deviation range. We keep formalising this argument more carefully for future work. However, in an experiment conducted as part of this study that takes two identical complete graphs and attempts to methodically remove one edge at random without replacement, varying the number of edges and the number of nodes in the graph simultaneously yields the results given in Fig.˜2. Interestingly, as edges and nodes are reduced, the energy decreases up to edge number 6, but then rises again for sparser graphs. This non-monotonic trend may reflect finite-size effects, ansatz expressibility limits, or optimizer artifacts, and warrants further investigation and study.
((a))
((b))
Figure 2: The illustration of variation in QAOA energies with simultaneous edge and node reduction. (a) Shows how edges were methodically removed from a five node complete graph (the graphs are reduced from (j) $\rightarrow$ (a)). A copy of each graph was used when testing, thereby yielding a pair of identical graphs, on which the graph isomorphism formulation was carried out. (b) Shows the QAOA energies of the Hamiltonian for each of these graphs for which the isomorphic pairs has the quadratic program yield a function value of zero
V.2 VQE Results
The findings obtained from QAOA were further corroborated by the results of the VQE. In particular, VQE yielded an optimal function value of zero for isomorphic graph pairs and a non-zero value for non-isomorphic pairs, in agreement with the theoretical expectations of the QUBO formulation. This outcome demonstrates that the variational principle underlying VQE is able to faithfully capture the structural equivalence encoded in the Hamiltonian, thereby validating the observations made with QAOA. A notable feature of the VQE results is that its ground-state energies were of much larger magnitude than those from QAOA. This stems from VQE’s deeper ansatz and broader variational search, which approximate the ground state with higher fidelity.
Although the absolute energy scales differ, both methods showed consistent trends: isomorphic graph families clustered tightly in energy, while non-isomorphic pairs deviated. This clustering thus emerges as a variational signature of graph isomorphism and supports the use of energy landscapes as a heuristic for distinguishing graph families. It is important to emphasize, however, that this improved fidelity comes at a substantial computational cost. Unlike QAOA, which requires only a shallow circuit and a limited parameter set, VQE relies on repeated circuit preparation and measurement at each optimization step. In our implementation, this involved up to 600 iterations of the SPSA optimizer on a TwoLocal ansatz, resulting in significantly longer execution times and higher resource usage. This trade-off highlights a key distinction between the two approaches: QAOA offers a lightweight variational heuristic capable of rapidly identifying clustering patterns, whereas VQE provides a more resource-intensive but rigorous validation of those patterns. The convergence of VQE energies across iterations for multiple runs on four-node graph pairs is shown in Fig.˜3.
Taken together, these results suggest a complementary role for the two algorithms. QAOA, with its shallow-depth circuits, can serve as a practical diagnostic tool for detecting candidate isomorphic families, while VQE, through Ritz’s variational principle, can confirm these findings. The fact that both methods independently reproduce the clustering phenomenon strengthens the central claim of this work: that variational algorithms provide a viable pathway toward quantum heuristics for the Graph Isomorphism problem, even in the small-instance regime accessible on near-term devices. Furthermore, the consistency between QAOA and VQE supports the broader hypothesis that energy clustering is not an artifact of a particular ansatz or optimization method, but rather an intrinsic property of the Hamiltonian encoding of isomorphism.
((a))
((b))
Figure 3: The illustration of the energy for VQE (y-axis) for pairs of graphs with four nodes in respect with number of iterations (x-axis). The colors represent different runs on a class of graphs. It can be seen that after 400400 iterations the energies cluster.
VI Conclusion and Future Work
We have tested two different quantum algorithmic approaches, namely the QAOA and VQE, towards obtaining solutions of small case-instances for the Graph Isomorphism Problem, where the Hamiltonian is formulated using the QUBO formulation. We have shown through numerical simulations that the results of both algorithms support each other in identifying isomorphic and non-isomorphic pairs for four, and five nodes in a non-noisy device. Thus, we have proposed a new application area for variational quantum algorithms.
We have demonstrated some empirical evidence that agree with previous literature Szegedy (2019) in regards to QAOA energies of isomorphic graphs clustering together. Additionally, we have also shown that similar phenomena happen for VQE simulations as well. A further extension of the investigation pursued in this work involves expanding the idea of energies clustering for families of isomorphic graphs and running the formulation on current devices with greater qubit number. Another natural next step for this line of research is to scale the variational formulations to graphs with 6–8 nodes using more efficient encodings that reduce qubit overhead. While the quadratic growth of qubits in the QUBO formulation presents a bottleneck, recent work on compact encodings and symmetry exploitation could make moderately larger instances tractable. In parallel, testing these formulations on real quantum hardware such as IBM’s superconducting processors or Google’s Sycamore device would provide valuable insights into the effects of noise, decoherence, and limited circuit depth. Employing error mitigation strategies, including zero-noise extrapolation, measurement error correction, and dynamical decoupling, would allow us to benchmark the robustness of the observed energy clustering under realistic device conditions and thereby bridge the gap between proof-of-principle simulations and experimental feasibility.
Acknowledgements.
We thank Dr. Farhan T. Chowdhury at the Ohio State University for a thorough evaluation of our initial draft and advising on key changes that we have implemented in the final manuscript. TC and SIM would like to thank Dr. Omar Shehab from IBM Quantum for insightful discussions during the sessions of Banglay Quantum Computing, and Dr. Amit Saha for his unrelenting support. AK would like to thank Dr. Özlem Salehi for the fruitful discussion regarding the QUBO formulation, and acknowledges support from the National Science Centre (NCN), Poland, under project number 2019/33/B/ST6/02011.
Data Availability Statement
The data and codes that support the findings of this study are openly available at https://github.com/LegacYFTw/Graph-Isomorphism.